Solve for X Show algebraic steps.
4/x+3 - 2/x-3 = 2

Answer by bsmath(21)

4/x+3 - 2/x-3 = 2
First, multiply through by x to get the standard form of the Quadratic Formula.
4+3x-2-3x=2x, it looks like we do not have the Quadratic Formula form, it is
linear in x. All we have to do is combine terms and solve. 4-2=2 and 3x-3x=0,
so we have 2=2x, divide both sides by 2 gets us x=1.